A sealed container holding 0.0255 L0.0255 L of an ideal gas at 0.997 atm0.997 atm and 73 ∘C73 ∘C is placed into a refrigerator and cooled to 35 ∘C35 ∘Cwith no change in volume. Calculate the final pressure of the gas.
Answer #1

pressure P1 = 0.997 atm

pressure P2 = ??

temperature T1 = 73 oC = 346 K

T2 = 35 oC = 308 K

P1 / T1 = P2 / T2

0.997 / 346 = P2 / 308

P2 = 0.888

final pressure = 0.888 atm